mathslover (mathslover):

without multiplying show that 2*9 is greater than 1*10

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (kainui):

Show that 10>18 ok np

mathslover (mathslover):

oops sorry wait

mathslover (mathslover):

edited

mathslover (mathslover):

sorry for the mistake

OpenStudy (lgbasallote):

without multiplying...then add

OpenStudy (lgbasallote):

9+ 9 > 10

mathslover (mathslover):

haah good @lgbasallote

OpenStudy (anonymous):

divide both sides by 2 1*9 and other side 1*5 now divide both sides by 1 9>5 hence proved

OpenStudy (kainui):

My logic was really messed up on this one honestly. I was thinking, well, anything multiplied by 1 is really just that number, so I'm only proving 2*9 is greater than 10. Then I thought, well, 9 is only 1 less than 10, and any number greater than 1 that's one less than another number who's a multiple has to be larger than that other number.

OpenStudy (shubhamsrg):

how about this fun ques: which is greater ? e^(pi) or pi^e ??

OpenStudy (kainui):

Is there a way of thinking about this without just directly using a calculator? I kind of approximated e=2 and pi=3 and then figured 8<9, so thought pi^e must be larger, but its not. That approximation isn't really right.

OpenStudy (shubhamsrg):

why you took e=2 when actually e=2.7something ..

OpenStudy (kainui):

Just because I could do that in my mind. As long as it kept e<pi I figured it'd be fine. I can easily do 2^3 and 3^2 in my head lol.

mathslover (mathslover):

2^100 > 100^2

mathslover (mathslover):

this is the logic

OpenStudy (anonymous):

see this \[\log e ^{\pi}= \pi \log e=\pi\] \[\log \pi ^{e} = e \log \pi =e \log (22/7) = e ( \log 22 -\log 7)\]

mathslover (mathslover):

though i know that: \[\large{e^{i\pi}=-1}\]

OpenStudy (shubhamsrg):

but i want it without much use of calculator..

OpenStudy (shubhamsrg):

@theyatin

mathslover (mathslover):

yes @shubhamsrg it is simple.. e = 2 (approx.) pi = 3(approx.) 2^3 < 3^2

OpenStudy (anonymous):

then do it as maths lover is doing :P

mathslover (mathslover):

:D.. logicana

OpenStudy (shubhamsrg):

well your ans is wrong..

OpenStudy (anonymous):